OpenStudy (anonymous):

When graphing the following equations, which one will be a straight line? A. y = x^2 + 2 B. f(x) = x + 1 C. y = |x + 4| D. x2 + y^2 = 1

OpenStudy (anonymous):

Help plz :/

OpenStudy (anonymous):

Well, i think it would be B.

OpenStudy (anonymous):

@UserA113 Do you have to use Pythagorean Theorm? or another one o-o

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

gtfo @userA113 lol what do you think your doing ?

OpenStudy (anonymous):

OOO. How fun. xD Urm, Well you know the a^2+b^2=c^2. All you have to do is plug in the numbers.

OpenStudy (anonymous):

sorry mate @eddielikestacos

OpenStudy (anonymous):

i still need help lol

OpenStudy (anonymous):

In A and D, you values that are increasing by multiplication. They will end up growing at a faster rate than numbers that are added. So the answer is B and C because they increase at a constant rate. Does that make sense?

OpenStudy (anonymous):

Well, When you Graph a Variable with an exponent, it will curve. and obviously you dont want a graph that curves. So A and D is Gone. C is an Absolute value which will give you 2 Graphs, so that is gone. so the only one would be B because if you graph y=x+1, Its a straight line. if any of this made sense xD